Calculations Saved P12-10 Calculating Real Returns and Risk Premiums (LO1) Youve observed the following nominal returns on
Answer #1

1. Average return = (0.05 – 0.11 + 0.29 + 0.16 + 0.12) / 5 = 0.102 or 10.2%

(1 + R) = (1 + r)(1 + h)

r = (1.102/1.038) – 1

r = 0.0617 or 6.17%

2. Average risk premium = Average return − Average risk-free rate

Average risk premium = 0.102 − 0.0525

Average risk premium = 0.0495 or 4.95%